Laura has $4.50 in dimes and quarters. She has 3 more dimes than quarters. How many quarters does she have?

Answer by checkley79(3341)

D=Q+3
.10D+.25Q=4.50
.10(Q+3)+.25Q=4.50
.10Q+.30+.25Q=4.50
.35Q=4.50-.30
.35Q=4.20
Q=4.20/.35
Q=12 QUARTERS.
D=12+3=15 DIMES.
PROOF:
.10*15+.25*12=4.50
1.50+3.00=$4.50